A business man loaned P500,000 from a local bank that charges an interest rate of 12%.

A. how much is he going to pay at the end of 5 years if it was made at 12% simple interest.

B. how much is he going to pay at the end of 6 years if it was made at 12% compounded quarterly.

C. how much is he going to pay at the end of 7 years if it was made at 12% compounded continuously.

-this is an continuous compounding

with solution
Answer #1

A.Amount at the end of 5 years if it was made at 12% simple interest.

=500,000+ [500,000*12%*5 years]

= 800,000

B. Amount at the end of 6 years if it was made at 12% compounded quarterly.

=500,000(1+12%/4)6*4

=500,000(1.03)24

=500,000*2.03279410646

=1016397.05

C.Amount at the end of 7 years if it was made at 12% compounded continuously

=P*ei*t   e=2.7183 i=12%or 0.12 t=7 years

=500,000*(2.71830.12*7)   

=500,000* 6.2965

=31,48,250